OpenStudy (highschoolmom2010):

You roll two standard number cubes. What is the probability that the sum is even, given that one number cube shows a 2?

OpenStudy (anonymous):

did you list all the possibilities pairs that has 2?

OpenStudy (highschoolmom2010):

no....

OpenStudy (anonymous):

well are you going to?

OpenStudy (highschoolmom2010):

1,1 1,2 1,3 1,4 1,5 1,6 2,1 2,2 2,3 2,4 2,5 2,6 3,1 3,2 3,3 3,4 3,5 3,6 4,1 4,2 4,3 4,4 4,5 4,6 5,1 5,2 5,3 5,4 5,5 5,6 6,1 6,2 6,3 6,4 6,5 6,6 so probability of at least one cube shows a 2 is 11/36.

jimthompson5910 (jim_thompson5910):

It says "given that one number cube shows a 2", so let's assume that it's the first number cube that shows a 2. If that's the case, then we look at row 2 which is 2,1 2,2 2,3 2,4 2,5 2,6 now what are the sums of each pair?

OpenStudy (highschoolmom2010):

3,4,5,6,7,8

jimthompson5910 (jim_thompson5910):

which of those sums are even?

OpenStudy (anonymous):

all possible pairs that has 2 is: (1,2) (2,1) (2,2) (2,3) (3,2) (2,4) (4,2) (2,5) (5,2) (2,6) (6,2) So there are 11. Out of those 11 pairs, which one has the sum of even?

OpenStudy (highschoolmom2010):

(2,2) (2,4) (4,2) (2,6) (6,2)

OpenStudy (anonymous):

so that would be 5/11

OpenStudy (highschoolmom2010):

my book says 1/2.......

OpenStudy (highschoolmom2010):

but i dont see how

jimthompson5910 (jim_thompson5910):

If the first die is 2, then we have this set of possibilities 2,1 2,2 2,3 2,4 2,5 2,6 They add to: 3, 4, 5, 6, 7, 8 Of that list, only 4, 6, 8 are even. This is 3 sums out of 6 so, 3/6 = 1/2 is your answer

OpenStudy (anonymous):

Then the question assume only ONE of the cube shows 2. Show the pair (2,2) wasn't not supposed to be counted. So answer is indeed 5/10 = 1/2

OpenStudy (anonymous):

But really, the book wasn't being specific because such phrase "given that one number cube shows a 2" actually means AT LEAST one.

OpenStudy (highschoolmom2010):

so i only needed 1 die with a 2 on it not both of them

OpenStudy (anonymous):

Yes, in this case. Sometimes the misuse of language can change the meaning of a question. But as long as it's not a conceptual misunderstanding, then you'll be fine.

OpenStudy (highschoolmom2010):

ty @jim_thompson5910 @@sourwing